SLAC PEP-II
BABAR
SLAC<->RAL
Babar logo
HEPIC E,S & H Databases PDG HEP preprints
Organization Detector Computing Physics Documentation
Personnel Glossary Sitemap Search Hypernews
Unwrap page!
Det. Search
Who's who?
Meetings
FAQ
Images
Archive
Systems
Performance
Intern. region
Vertex Tracker
Drift chamber
DIRC
Calorimeter
IFR
LST
Magnet
Electronics
Trigger
Operations
Run Coordination
Contact Experts
Shift Takers Info
Operations Manual
Electronic Logbook
Ops Hypernews
Shift Signup
Check this page for HTML 4.01 Transitional compliance with the
W3C Validator
(More checks...)

Minutes of the Trigger Operation Meeting (8th May, 2007)

Attendees: Debbie Bard, Rainer Bartoldus, Al Eisner, Fang Fang, Kim Hojeong, Arik Kreisel, Selina Li, Rahmat Rahmat, Su Dong.

Operations

  • Shutdown to replace Q2/Q4 magnets is on-going. Two EMC modules (forward bottom) were switched off over teh weekend to let them dry out, and the effect was seen in both the EMT and TSF plots. We should check this is really the cause by recreating the plots with DCT-only and EMT-only triggers, perhaps as default in the L1T archive.
  • Al noticed that the L1 COSMICS trigger mask is not the same as the L1 COSMICS_L1PASSTHRU mask (see hypernews). COSMICS_L1PASSTHRU is as expected from TrgConfig, but COSMICS is different. Was something changed by hand and never changed back? Investigation after the meeting shows the default mask is correct - it seems experts are not returning to default after doing a custom mask. See followup to above hypernews.
  • In any case, COSMICS_L1PASSTHRU should not be the default cosmics configuration, as currently stated in the care and feeding manual. Debbie changed this after the meeting.
  • As Hojeong is the second contact for the trigger group, she should get a pager, or one of those new-fangled cell phones.

EMT

  • Once the EMC is turned on fully (and finished with calibration issues), we should testsome spare TPBs (carefully!).
  • A new version of L1EmtSim (based on bukind-20070122 tag) is in the nightlies.

Online

  • Want to re-test GLT libraries when the detector is up again (pref. with magnet). Request put on whiteboard but there may be very limited time when the magnet is up without beams!
  • Memory test for 3 boards at once is working, still working on TSF-ZPD calibration.

L3

  • See hypernews posting - question about exact meaning of turning L3OutDCH on and off. L3OutDCH also contains BhaBha veto, so it would be better to use a script line veto if you want to remove tracks.

Data Quality and Monitoring

  • Data so far looks fine.
  • Something changed in the code to cause the TRG stripcharts to break, loosing all stripcharts follwing TRG in the file.

This page is created by Debbie Bard (djbard@slac.stanford.edu); however, you are welcome to edit and/or correct above items, specially for those who attended the meeting. Thanks.